OpenStudy (kity_kisses1234):

A certain forest covers an area of 2200km^2. Suppose that each year this area decreases by 3.75%. What will the area be after 7years? round your answer to the nearest square kilometer.

OpenStudy (kity_kisses1234):

@ganeshie8 please help me

OpenStudy (mrnood):

Suppose it decreased by 1% each year then you can see that each year there remains 0.99 times the original. What is the the equavalent proportion using 3.75%?

OpenStudy (kity_kisses1234):

um im not sure to be honest... math is really hard for me ive tried doing to problem myself like 20 times but get wrong each time

OpenStudy (mrnood):

I've given you the method above if it was 1% the proposrtion left is 1 - 1/100 = .99 (each year) if it was 2% then the proportion left is 1 - 2/100 = .98 (each year) If it was 3.5% then the the proprtion left would be 1 - 3.5/100 = .965 (each year) so what would the number be in your question when 3.75% is lost?

OpenStudy (kity_kisses1234):

.9625?

OpenStudy (mrnood):

Yes - good now- that amounr is left after each year so if you start with 2200 sq km then the table shows what happens year by year:  so what happemns after 7 years?